Part A How does the magnetic field due to a long straight wire carrying current vary with distance from the wire, ? It is constant. o 1/2 1/r Submit Request Answer

How does the magnetic field due to a long straight wire carrying current vary with distance from the wire, ? 
Answer #1

a) The magnetic field due to long straight wire is,

$$ B=\frac{\mu_{o} I}{2 \pi r} $$

That is magnetic field is inversely proportional to distance from the wire

$$ B \alpha \frac{1}{r} $$

That is the correct option is last option